# Meta-Learning Neural Process for Implied Volatility Surfaces with SABR-induced Priors

### v1

@@ 4.2 Overall and Stratified Performance
Table 2: Overall and stratified performance comparison. Short-, mid-, and long-term correspond to maturities of 0-3 months, 3-12 months, and >12 months, respectively. All errors are in basis points (BPS). Best result in each category is in bold.
@@
RMSE MAE RMSE MAE RMSE MAE RMSE MAE
VolNP-FT 99.60 48.86 106.82 51.40 77.06 40.60 97.17 58.39
VolNP-Base 166.85 49.77 174.55 46.01 150.07 56.09 121.79 77.66
SABR 153.71 67.59 135.45 59.83 158.84 70.95 338.48 194.22
GP 274.90 73.84 58.53 32.62 255.10 78.69 1244.98 838.73
SSVI 248.46 174.26 219.45 155.14 260.27 192.15 531.22 417.38
Table 2 presents the main results. Overall, VolNP-FT achieves the best performance, with the lowest RMSE (99.60 BPS) and MAE (48.86 BPS). A key finding comes from the comparison with VolNP-Base. While their overall MAE is similar, the pre-training strategy allows VolNP-FT to reduce the RMSE by nearly 40%. This suggests that the SABR-induced prior does not just improve the average fit, but is particularly effective at preventing large, financially significant prediction errors.
The stratified results provide further insight. For short-term options, where market data is most dense, the flexible Gaussian Process model performs best. However, as maturity increases and data becomes sparser, the performance of GP and other benchmarks degrades. In these mid- and long-term regimes, VolNP-FT outperforms all other models. This indicates that the learned geometric prior is crucial for constructing a coherent and robust surface, especially in regions with fewer data points.
Figure 2 reveals the models’ error patterns. Parametric models like SABR and SSVI show systematic errors on the ’wings’ of the surface, whereas VolNP-FT demonstrates uniformly low error, indicating superior generalization. The comparison with VolNP-Base highlights that the SABR-based prior acts as a powerful regularizer, suppressing large errors and ensuring a more robust surface reconstruction.
@@ 4.3 Sensitivity to Data Sparsity
To study the robustness of the model, we examine the performance of model under variety data sparsity (Figure 3). The RMSE metric shows that VolNP-FT maintains a lower RMSE across all levels of sparsity, demonstrating data efficiency. VolNP-FT’s lower RMSE confirms its ability to maintain a globally plausible surface shape instead of just fitting local points. This highlights its reliability for practical applications.
Figure 3: Model error (RMSE and MAE) as a function of the number of context points. VolNP-FT demonstrates superior data efficiency, achieving lower error with fewer context points and showing more stable performance as data becomes sparse.
@@ 4.4 Case Study: Surface Reconstruction Under Market Stress
We visualize the reconstructed surfaces on a challenging day: March 10, 2023, a day of significant market turmoil following the collapse of Silicon Valley Bank. This event provides a real-world stress test for model robustness. Figure 4 provides a comprehensive comparison, where each model reconstructs the surface from the same 100 context points drawn from the sparse market data (plotted as blue dots in panel a).Red dots in (a) are unobserved data to represent the ground truth.
@@
The benchmark models exhibit distinct failure modes. The SABR model (b) produce a surface that is structurally plausible but overly rigid, failing to capture finer market details. The highly flexible Gaussian Process (c) overfits the sparse data, producing an unstable and financially implausible surface with unnatural oscillations. The ablation model VolNP-Base (e), while capturing some features of smiles, yields an irregular surface, suggesting it overfits local noise without a global structural guide. In contrast, VolNP-FT (f) reconstructs a surface that is successfully capturing the essential features of the ground truth market data. This visual evidence reinforces the quantitative results, suggesting that the geometric prior learned during pre-training is critical for regularizing the model and enabling it to generate coherent global structures from sparse local information.
@@
Figure 4: Implied volatility surface construction on March 10, 2023. Each model attempts to reconstruct the full surface from the same context dots (blue dots) shown in (a). Red dots in (a) are unobserved data to represent the ground truth.
@@
A critical test for a model’s practical utility is the absence of butterfly arbitrage. We test for violations of the no-butterfly-arbitrage condition, which ensures the model’s implied risk-neutral probability density is non-negative. This test is performed by checking the Durrleman condition (further details, see Theorem 2.9, condition IV3 of (Roper, 2010)). Figure 5 shows the result on March 10, 2023. Smiles from VolNP-Base (bottom row) exhibit significant violations of this condition, indicated by the red shaded areas where arbitrage opportunities exist. This makes the model unusable in practice. In contrast, smiles from VolNP-FT (top row) are consistently free of such arbitrage. This shows that our two-stage training successfully provide a structural prior that guides the model to produce outputs that adhere to fundamental financial principles.
@@
Figure 5: Arbitrage analysis on March 10, 2023. The red shaded areas indicate regions where the smile is admitting butterfly arbitrage (against Durrleman condition).
@@ 5 Conclusion
Traditional methods for constructing the implied volatility surface often treat each trading day as an isolated calibration task. This paper explores an alternative approach by framing surface construction as a meta-learning task. We propose the Volatility Neural Process, a model designed to learn a general mapping from sparse option quotes to a complete implied volatility surface. The model is first pre-trained on SABR-generated data to learn a financial-theoretic prior, and subsequently fine-tuned on market data to capture real-world dynamics. By learning the general task of surface construction, this framework yields a single model that does not require daily re-calibration, offering a potentially more efficient and reliable solution for constructing the implied volatility surface.
Empirical results on the SPX options dataset indicate that this approach is effective. The proposed model outperforms the considered benchmarks in overall accuracy and demonstrates robustness to data sparsity. The analysis also confirms that the pre-training stage is crucial for reducing large errors and for generating surfaces that are free from butterfly arbitrage.

### v3

@@ 4.2 Overall and Stratified Performance
Table 2: Overall and stratified performance comparison. Short-, mid-, and long-term correspond to maturities of 0-3 months, 3-12 months, and >12 months, respectively. RMSE and MAE are in basis points (BPS). Best result in each category is in bold.
@@
RMSE MAE R2 RMSE MAE R2 RMSE MAE R2 RMSE MAE R2
VolNP-SP 96.04 48.22 0.98 104.03 50.44 0.98 71.80 41.48 0.99 85.44 53.33 0.98
VolNP-Base 164.75 48.98 0.96 171.29 43.72 0.96 151.41 59.01 0.96 121.06 79.14 0.95
SABR 153.71 67.59 0.96 135.45 59.83 0.97 158.84 70.95 0.95 338.48 194.22 0.63
GP 274.90 73.84 0.88 58.53 32.62 0.99 255.10 78.69 0.88 1244.98 838.73 -3.94
SSVI 268.39 151.34 0.88 253.84 143.78 0.90 269.31 144.51 0.86 464.59 346.66 0.31
Table 2 presents the errors across all trading days. Overall, VolNP-SP achieves the best performance, with the lowest RMSE (96.04 BPS) and MAE (48.22 BPS), and the highest R2 (0.98). A key finding comes from the comparison with VolNP-Base. While their overall MAE is similar, the pre-training strategy allows VolNP-SP to reduce the RMSE by nearly 40%. This suggests that the SABR-induced prior does not just improve the average fit, but is particularly effective at preventing large, financially significant prediction errors.
The stratified results provide further insight. For short-term options, where market data is most dense, the flexible Gaussian process model performs best. However, as maturity increases and data becomes sparser, the performance of GP and other benchmarks degrades. In these mid- and long-term regimes, VolNP-SP outperforms all other models. This indicates that the learned geometric prior is crucial for constructing a coherent and robust surface, especially in regions with fewer data points.
We complement the quantitative metrics with two heatmaps that aggregate errors over the full $(k,\tau)$ domain across all test days. For each test day we bin target coordinates $(k,\tau)$ on a fixed global grid. Within each bin we compute the per-day RMSE and the per-day mean signed error (bias). We then take the median across test days in each bin.
The RMSE heatmaps (Figure 3) confirm that VolNP-SP maintains low error not only near-the-money but also across sparse long-maturity regions where benchmarks perform poorly. The bias heatmaps (Figure 4) show SABR and SSVI display pronounced directional biases on wings or long maturities, while VolNP-SP’s bias is near-neutral and spatially smooth, consistent with its higher overall R2 and lower RMSE.
@@
Figure 4: Mean signed error (bias) heatmaps for benchmark and VolNP models. Blue indicates underestimation; red indicates overestimation.
4.3 Monthly RMSE Time Series
To examine temporal robustness, we aggregate daily RMSE into monthly averages. This view highlights regime shifts and macro stress episodes, and contrasts the stability of VolNP-SP against the market-only baseline.
Figure 5: Monthly averaged RMSE (in BPS) over time, stratified by maturity. Each subplot compares VolNP-SP (blue) and VolNP-Base (orange).
Figure 5 reveals three patterns. First, during the early-2020 volatility spike, both models suffer higher errors, but VolNP-SP rises less across all maturities. Second, from 2021 onward, VolNP-SP maintains a persistent gap over VolNP-Base in mid-term and long-term, aligning with its superior in stratified metrics. Third, short-dated months exhibit small fluctuations where both models are close, but VolNP-SP remains at least as accurate.
4.4 Computational Efficiency
We report the average runtime per-day evaluation [^5] Environment. Python 3.10; Intel i7-12700 (2.10 GHz); NVIDIA RTX 3060 (12 GB). VolNP models run in PyTorch (CUDA) and we time a single forward pass; SABR/SSVI/GP run on CPU (NumPy/SciPy; scikit-learn for GP) and include calibration/fit plus prediction.. One evaluation corresponds to selecting $N=100$ context quotes for a test day and producing predictions for all remaining target quotes of that day in a single run. Table 3 lists the mean time per evaluation in milliseconds.
Table 3: Average per-day evaluation runtime (milliseconds).
VolNP-Base VolNP-SP SSVI GP SABR
Avg time (ms) 23.9 24.5 123.4 189.6 1211.1
Both VolNP variants require about 25 ms per evaluation, and outperform other methods. This reflects deployment cost for generating a day’s IVS from sparse context, and VolNP scales well for intraday updates.
@@ 4.5 Sensitivity to Data Sparsity
To study the robustness of the model, we examine the performance of the model under variety data sparsity (Figure 6). The RMSE metric shows that VolNP-SP maintains a lower RMSE across all levels of sparsity, demonstrating data efficiency. VolNP-SP’s lower RMSE confirms its ability to maintain a globally plausible surface shape instead of just fitting local points. This highlights its reliability for practical applications.
Figure 6: Model error (RMSE and MAE) as a function of the number of context points. VolNP-SP demonstrates superior data efficiency, achieving lower error with fewer context points and showing more stable performance as data becomes sparse.
@@ 4.6 Case Study: Surface Reconstruction Under Market Stress
We visualize the reconstructed surfaces on a challenging day: March 9, 2020, during the COVID-19 volatility shock. This event provides a real-world stress test for model robustness. Figure 7 provides a comprehensive comparison, where each model reconstructs the surface from the same 100 context points drawn from the market data (plotted as blue dots in panel a). Red dots in (a) are unobserved data to represent the ground truth.
@@
The benchmark models exhibit distinct failure modes. The SSVI (d) and SABR (b) shows clear volatility smiles but are overly rigid, failing to capture finer market details. The highly flexible Gaussian process (c) overfits the sparse data, producing an unstable and financially implausible surface with unnatural oscillations. The ablation model VolNP-Base (e), while capturing some features of smiles, yields an irregular surface, suggesting it overfits local noise without a global structural guide. In contrast, VolNP-SP (f) reconstructs a surface that successfully captures the essential features of the ground truth market data. This visual evidence reinforces the quantitative results, suggesting that the geometric prior learned during pre-training is critical for regularizing the model and enabling it to generate coherent IVS from sparse dataset.
@@
Figure 7: Implied volatility surface construction on March 9, 2020. Each model attempts to reconstruct the full surface from the same context dots (blue dots) shown in (a). Red dots in (a) are unobserved data to represent the ground truth.
@@
We test for violations of the no-butterfly-arbitrage condition, which ensures the model’s implied risk-neutral probability density is non-negative. This test is performed by checking the Durrleman condition (further details, cf. Theorem 2.9, condition IV3 of (Roper, 2010)). Figure 8 shows the result on March 9, 2020. Smiles from VolNP-Base (bottom row) exhibit significant violations of this condition, indicated by the red shaded areas where arbitrage opportunities exist. This makes the model unusable in practice. In contrast, smiles from VolNP-SP (top row) are consistently free of such arbitrage. This shows that our two-stage training successfully provides a structural prior that guides the model to produce outputs that adhere to fundamental financial principles.
@@
Figure 8: Arbitrage analysis on March 9, 2020. The red shaded areas indicate regions where the smile admits butterfly arbitrage (against Durrleman condition).
We further visualize two arbitrage diagnostics. First, for fixed maturities $\tau\in\{0.50,1.00,2.50\}$ we evaluate Black-Scholes call prices $C$ on a uniform strike grid around the forward and compute the discrete second difference $B(K)=C(K+h,\tau)-2C(K,\tau)+C(K-h,\tau)$ as a proxy for convexity in $K$. Nonnegative $B(K)$ rules out butterfly arbitrage. Figure 9 overlays the resulting curves. VolNP-SP yields smooth and largely nonnegative spreads across strikes, while the VolNP-Base shows negative values and oscillations on the wings, particularly at longer maturities.
Second, we examine the total variance $w(k,\tau)=\sigma^{2}(k,\tau)\,\tau$. In a calendar-arbitrage-free surface, $w(k,\tau)$ should rise with $\tau$ at fixed log-moneyness $k$ and vary smoothly across $k$. Figure 10 shows families of $w(\cdot,\tau)$ curves. VolNP-SP exhibits nested, coherent term structures with well-behaved wings. VolNP-Base develops local humps at extreme $k$ that may compromise monotonicity. These views highlight the regularizing role of the SABR-induced prior under market stress.
Figure 9: Butterfly-spread diagnostics on March 9, 2020. For $\tau=0.50$, $1.00$ and $2.50$, we compute discrete second differences of call prices on a uniform strike grid using each model’s IV. Green curve is the result of VolNP-SP. Orange curve is the result of VolNP-Base.
Figure 10: Total variance $w(k,\tau)$ on March 9, 2020. Curves show $w(\cdot,\tau)$ over log-moneyness for maturities from 0.50 year to 3.00 year.
@@ 5 Conclusion
This paper explores an alternative approach by framing surface construction as a meta-learning task. We propose the Volatility Neural Process, a model designed to learn a general mapping from sparse option quotes to a complete implied volatility surface. The model is first pre-trained on SABR-generated data to learn a financial-theoretic prior, and subsequently fine-tuned on market data to capture real-world dynamics. By learning the general task of surface construction, this framework yields a single model that does not require daily re-calibration, offering a potentially more efficient and reliable solution for constructing the implied volatility surface.
Empirical results on the SPX options dataset indicate that this approach is effective. The proposed model outperforms the considered benchmarks in overall accuracy and demonstrates robustness to data sparsity. The analysis also confirms that the pre-training stage is crucial for reducing large errors.
Future work. The prior should be viewed as a modeling choice rather than a fixed specification. SABR is merely one example. Alternative priors can be derived from structural models (e.g., Heston, rough volatility) via synthetic surface generation, or from arbitrage-free parameterizations (e.g., SSVI). Future work can compare and combine priors through mixture or ensemble schemes and develop regime-adaptive procedures for prior selection.